OpenStudy (bbb911):

This is Geometry: Find the area of a regular octagon with an apothem 13.3 feet long and each side 11 feet long.

OpenStudy (anonymous):

h=2*13.3=26.6 A=(11*26.6)/2=146.3 octagon have has 8 side it mean 8 triangle than time 8 A = 146.3*8 A=1,170,4 ft^2

OpenStudy (bbb911):

here is a multiple choice to that question A. 585.2 ft sq B. 368.9 sq C. 292.6 ft sq D. 73.15 ft sq

OpenStudy (bbb911):

Thank you A it is :D